# The $\ell$-test: leveraging sparsity in the Gaussian linear model for improved inference

### v1

@@ 3 $\ell$-test confidence intervals
[…]
@@
Computationally, one may be concerned that computing $\hat{C}_{j}$ requires many LASSO runs for a fine grid of $\gamma$ values. It is known (Efron et al., 2004, Rosset and Zhu, 2007) that the LASSO solution is piecewise linear in $\lambda$ and that these paths can be generated by efficient algorithms, but in Appendix D, we show that the LASSO solution ($\hat{\bm{\beta}}^{\lambda}_{-j}(\gamma)$) is also piecewise linear in $\gamma$ (for fixed $\lambda$) and we provide an algorithm to efficiently generate these paths as well. Combining these two path-generating algorithms (in $\lambda$ and $\gamma$) provides an efficient way to share computation to efficiently compute all the LASSO solutions needed for $\hat{C}_{j}$.
@@ 4 $\ell$-test inference conditional on LASSO selection
The $\ell$-test p-value’s distributional form (2.7) makes it extremely straightforward (both conceptually and computationally) to adjust it to be conditionally valid given $\hat{\beta}_{j}^{\lambda}\neq 0$: simply divide the $\ell$-test p-value $p_{j}^{\lambda}$ by $r_{j}^{\lambda}:=\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}\neq 0\mid\bm{S}^{(j)})=\mathbb{P}_{H_{j}}(u_{1}\notin[v_{-},v_{+}]\mid\bm{S}^{(j)})=1-F_{u}(v_{+})+F_{u}(v_{-})$ to get $\underline{p}_{j}^{\lambda}:=p_{j}^{\lambda}/r_{j}^{\lambda}$. The fact that, under $H_{j}$ $\underline{p}_{j}^{\lambda}$ has a $\mathrm{Unif}(0,1)$ distribution conditional on $\bm{S}^{(j)}$ and $\hat{\beta}_{j}^{\lambda}\neq 0$ follows because $r_{j}^{\lambda}$ is exactly the supremum value $p_{j}^{\lambda}$ can take as long as $\hat{\beta}_{j}^{\lambda}\neq 0$, and the density of $p_{j}^{\lambda}$ between 0 and $r_{j}^{\lambda}$ is uniformly distributed (see Section 2.3); it follows further that $\underline{p}_{j}^{\lambda}$’s null distribution conditional only on $\hat{\beta}_{j}^{\lambda}\neq 0$ is also $\mathrm{Unif}(0,1)$.
Corollary 4.1 (Validity of the conditional $\ell$-test).
@@
Now for obtaining post-LASSO-selection valid confidence interval for $\beta_{j}$, we need to invert a conditionally valid test for $H_{j}(\gamma),\gamma\in\mathbb{R}$. For testing $H_{j}(\gamma)$, following the suggestion in Section 3, the test statistic should be based on the LASSO estimate on $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$, that is $|\hat{\beta}_{j}^{\lambda}(\gamma)|$, whereas the model selection event is still based on the original, un-centered LASSO estimate, $\hat{\beta}_{j}^{\lambda}$. Furthermore, one can choose to use a different penalty parameter for the test statistic than the one used for the selection event. This prompts us to understand tests for $H_{j}(\gamma)$ based on $|\hat{\beta}_{j}^{\lambda_{\ell}}(\gamma)|$, valid conditionally on $\{\hat{\beta}_{j}^{\lambda_{s}}\neq 0\}$, where $\lambda_{\ell}$ and $\lambda_{s}$ need not be the same. Because now the test-statistic and the selection event are based on different LASSO estimates, a conditional p-value would not have such a simple form as for $\underline{p}^{\lambda}_{j}$, but we can still obtain valid p-value using CDF transforms if we can characterize the distribution of $|\hat{\beta}_{j}^{\lambda_{\ell}}(\gamma)|\mid\{\hat{\beta}_{j}^{\lambda_{s}}\neq 0\}$, under $H_{j}(\gamma)$.
First note that, as discussed in Section 3, Lemma 2.1 can be applied to $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$ under $H_{j}(\gamma)$ to show that $\bm{S}^{(j)}(\gamma)=(\bm{X}_{j}^{T}\bm{y},\hat{\sigma}_{j}(\gamma))$ is sufficient under $H_{j}(\gamma)$, where, $\hat{\sigma}_{j}(\gamma)=\|(\bm{I}-\bm{P}_{-j})(\bm{y}-\gamma\bm{X}_{j})\|$, and that $\bm{y}-\gamma\bm{X}_{j}$ can be written as $\bm{P}_{-j}\left(\bm{y}-\gamma\bm{X}_{j}\right)+\hat{\sigma}_{j}(\gamma)\bm{V}\bm{u}^{\gamma}$, with $\bm{u}^{\gamma}\mid\bm{S}^{(j)}(\gamma)\stackrel{{\scriptstyle H_{j}(\gamma)}}{{\sim}}\mathrm{Unif}(\mathbb{S}^{n-d})$. In light of this result, one can now apply Theorem 2.1 to $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$ to conclude that $\hat{\beta}_{j}^{\lambda_{\ell}}(\gamma)\leq b$ if and only if $u_{1}^{\gamma}\leq\Lambda_{j}^{*}(b,\mathrm{sign}(b);\gamma)$, where $\Lambda_{j}^{*}(b,\epsilon;\gamma)$ is exactly equal to $\Lambda(b,\epsilon)$ but with $\bm{y}$ replaced with $\bm{y}-\gamma\bm{X}_{j}$ and $\mathrm{sign}(0):=1$. In fact Theorem A.1 (that characterizes the distribution of $\hat{\beta}_{j}^{\lambda}\mid\bm{S}^{(j)}(\gamma)$ under $H_{j}(\gamma)$) in Appendix A shows that this same $u_{1}^{\gamma}$ can be used to characterize the event $\{\hat{\beta}_{j}^{\lambda_{s}}=0\}$, stating it is equivalent to $\{u^{\gamma}_{1}\in[\Lambda_{j}(0,\pm 1;\gamma)]\}$, where,
@@
Like Liu et al. (2018) but unlike, e.g., Lee et al. (2016), our conditional inferences do not condition on anything about the LASSO’s selection except that it selects the $j^{\mathrm{th}}$ coefficient. As mentioned in Section 1.3, the key difference between our conditional inference and Liu et al. (2018)’s is essentially the same as the difference between the (unconditional) $\ell$-test and the $t$-test, and indeed in Section 5.4 we find that our conditional confidence intervals improve over those of Liu et al. (2018) similarly to how the unconditional $\ell$-test confidence intervals from Section 3 improve over standard $t$-test confidence intervals.

### v2

@@ 2.5 Putting it all together: the $\ell$-test
2.6 Asymptotic extension beyond the Gaussian linear model
Due to the close relationship between the Gaussian linear model and the multivariate Gaussian distribution, it is intuitively easy to extend the $\ell$-test to a large class of models that admit asymptotically Gaussian estimators. In particular, consider a sequence of models and for each model, a parameter vector $\bm{\theta}_{n}$ in a local asymptotic regime, i.e., $\bm{\theta}_{n}=\bm{\theta}/\sqrt{n}$ for some fixed vector $\bm{\theta}$. Then for any estimator $\hat{\bm{\theta}}_{n}$ such that $\sqrt{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{n})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(\bm{0},\bm{\Sigma})$ for some fixed positive definite covariate matrix $\bm{\Sigma}$ with its own consistent estimator $\hat{\bm{\Sigma}}_{n}$, we can compute response vector $\bm{y}=\sqrt{n}\hat{\bm{\Sigma}}_{n}^{-1/2}\hat{\bm{\theta}}_{n}$ and design matrix $\bm{X}=\hat{\bm{\Sigma}}_{n}^{-1/2}$ and apply the $\ell$-test to $(\bm{y},\bm{X})$, which asymptotically satisfies a Gaussian linear model with coefficient vector $\bm{\theta}$ and error variance $\sigma^{2}=1$ (see li2021whiteout for an asymptotic result like this for knockoffs). In fact, since $\sigma^{2}$ is known in this asymptotic regime, we can apply a version of the $\ell$-test that leverages the known $\sigma$ via some simple substitutions; for brevity of presentation, we defer these details to Appendix C. In the remainder of this section we will formalize this asymptotic application of the known-$\sigma$-$\ell$-test, but first pause to emphasize that this vastly expands the applicability of the $\ell$-test framework, e.g., to M-estimators (including the maximum likelihood estimator) in models that are differentiable in quadratic mean (vandervaart1998asymptotic). In particular, the ordinary least squares estimator paired with the robust sandwich variance estimator (white1980) allows the $\ell$-test to be applied asymptotically to a heteroskedastic linear model.
In addition to the aforementioned local asymptotic regime, we will need continuity and uniqueness of the cross-validation procedure in the $\ell$-test. These properties hold under very mild regularity conditions, but we defer their cumbersome statement to the appendix as Assumption 1.
Theorem 2 (The $\ell$-test for asymptotically Gaussian models).
For some $\bm{\theta}$ and $\bm{\Sigma}$, let $\bm{\theta}_{n}=\bm{\theta}/\sqrt{n}$ and suppose that the sequence of random variables $\left\{\left(\hat{\bm{\theta}}_{n},\hat{\bm{\Sigma}}_{n}\right)\right\}_{n}$ satisfies
$\sqrt{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{n})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(\bm{0},\bm{\Sigma})\qquad\textrm{ and }\qquad\hat{\bm{\Sigma}}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\bm{\Sigma}.$
Denote the known-$\sigma$-$\ell$-test $p$-value for testing $H_{j}$, applied to data $\left(\bm{y}=\sqrt{n}\hat{\bm{\Sigma}}_{n}^{-1/2}\hat{\bm{\theta}}_{n},\,\bm{X}=\hat{\bm{\Sigma}}_{n}^{-1/2}\right)$}, by $p_{j,n}$. Then under Assumption 1, $p_{j,n}$ is asymptotically valid for testing $\theta_{j}=0$, i.e., if $\theta_{j}=0$, then
$\underset{n\rightarrow\infty}{\lim\sup}~\mathbb{P}\left(p_{j,n}\leq\alpha\right)=\alpha,\;\forall\alpha\in[0,1].$
The proof of this theorem, available in Appendix D.6, relies on showing that the known-$\sigma$-$\ell$-test $p$-value is a continuous function of the data $(\bm{y},\bm{X})$, followed by an application of the continuous mapping theorem. As an intermediate result, the known-$\sigma$-$\ell$-test $p$-value using a fixed $\lambda$ is also asymptotically valid for the above class of models (without the need for Assumption 1), which we formally state in Corollary 3 of the appendix.
The understanding of the $\ell$-test’s power gain developed in Section 2.2 suggests that the asymptotic $\ell$-test described in Theorem 2 will outperform the standard asymptotic $z$-test when (1) $\bm{\theta}$ is sparse and (2) $\bm{\Sigma}$ has sufficient off-diagonal entries (the latter ensures the columns of $\bm{X}=\hat{\bm{\Sigma}}_{n}^{-1/2}\approx\bm{\Sigma}^{-1/2}$ are not too close to orthogonal).
3 Power analysis
As Section 2.2 argues, the power gain of the $\ell$-test under sparsity comes from the informative re-centering by $\hat{m}_{j}$ that guesses the sign of $\beta_{j}$. In this section we study this power gain theoretically by considering the test that simply re-centers $u_{1}$ by $\hat{m}_{j}$, i.e., it rejects for large values of $|u_{1}-\hat{m}_{j}|$ conditional on $\bm{S}^{(j)}$. This test gains power over the $t$-test via the same re-centering argument as the $\ell$-test does, but the latter is less tractable to analyze theoretically. To further ease our analysis, we treat $\sigma$ as known, allowing us to instead work with a known-$\sigma$-analogue of the statistic $|u_{1}-\hat{m}_{j}|$, details of which are discussed in Section C of the appendix. Finally, we treat $\lambda$ as fixed instead of chosen by cross-validation as in Section 2.4, and denote the $p$-value for this re-centered test when the sample size is $n$ as $p_{j,n}^{\mathrm{r},\lambda}$. We find in Appendix E.2 that this test (with $\lambda$ fixed at the mean of its cross-validated value computed from independent simulations) empirically performs indistinguishably from the (known-$\sigma$-analogue of the) $\ell$-test, suggesting that despite the simplifications we make in this section for mathematical tractability, both qualitative and quantitative conclusions from our power analysis should apply to the $\ell$-test as well.
The primary technical toolbox we rely on is Approximate Message Passing (AMP) (bayati2012), which allows us to precisely characterize the behavior of the LASSO estimate inside the known-$\sigma$-analogue of $\hat{m}_{j}$ under a proportional asymptotic limit. In particular, we assume the following random design and nuisance parameter regime.
Setting 1.
$d/n\to\kappa\in(0,1)$. The entries of $\bm{X}$ are drawn i.i.d. from $\mathcal{N}(0,1/n)$, the entries of $\bm{\beta}_{-j}$ are drawn i.i.d. from a distribution $B_{0}$ with finite second moment, and $\beta_{j}=h$ for some fixed scalar $h$ which we take, without loss of generality, to be non-negative.
This type of setting is standard in AMP theory, and will allow us to characterize the asymptotic $p$-values of the tests we are considering in terms of a particular bivariate Gaussian distribution: let
[…]
where $\alpha_{\lambda},\tau_{\lambda}$ satisfy the two AMP state-evolution equations (bayati2012) (using notation from crt_power as we use their results heavily in our proofs):
$\lambda=\alpha_{\lambda}\tau_{\lambda}\left(1-\kappa\mathbb{E}\left[\eta^{\prime}(B_{0}+\tau_{\lambda}W;\alpha_{\lambda}\tau_{\lambda})\right]\right);\hskip 19.91684pt\tau_{\lambda}^{2}=\sigma^{2}+\kappa\mathbb{E}\left[\left(\eta(B_{0}+\tau_{\lambda}W;\alpha_{\lambda}\tau_{\lambda})-B_{0}\right)^{2}\right],$
where $\eta(x;\omega):=\mathrm{sign}(x)(|x|-\omega)_{+}$ is the soft-thresholding function and $\eta^{\prime}$ denotes its derivative with respect to its first argument. Note that although the AMP state-evolution equations are nonlinear and do not admit a closed form solution for $(\alpha_{\lambda},\tau_{\lambda})$, they do uniquely determine them as a function of only $\sigma^{2}$, $\lambda$, $\kappa$, and $B_{0}$. Furthermore, $\tau_{\lambda}$ can be interpreted as a measure of the asymptotic error of the LASSO because the asymptotic mean squared estimation error of the LASSO is $\kappa^{-1}(\tau^{2}_{\lambda}-\sigma^{2})$ (bayati2012, Corollary 1.6).
First, $F$ represents the asymptotic distribution of the $z$-test statistic, so we can easily relate the one- and two-sided $z$-test $p$-values (the known-$\sigma$ analogues of the $t$-tests), denoted $p_{j,n}^{z,1}$ and $p_{j,n}^{z,2}$, respectively, to $F$ in Setting 1 (see Appendix D.7 for derivation):
$p_{j,n}^{z,1}\stackrel{{\scriptstyle d}}{{\longrightarrow}}1-\Phi\left(F\right)\quad\text{and}\quad p_{j,n}^{z,2}\stackrel{{\scriptstyle d}}{{\longrightarrow}}1-\left[\Phi\left(F\right)-\Phi\left(-F\right)\right]\mathrm{sign}(F),$
where the precise form of these expressions is chosen for ease of comparison with that for $p_{j,n}^{\mathrm{r},\lambda}$ in the following Theorem, whose proof appears in Appendix D.7.
Theorem 3.
Under Setting 1, $p_{j,n}^{\mathrm{r},\lambda}\stackrel{{\scriptstyle d}}{{\longrightarrow}}1-\left[\Phi\left(F\right)-\Phi\left(F-2G\right)\right]\mathrm{sign}(G).$
Comparing asymptotic expressions for $p_{j,n}^{z,1}$, $p_{j,n}^{z,2}$, and $p_{j,n}^{\mathrm{r},\lambda}$, we see that if $G=F$, then the re-centered test has the same asymptotic power as the two-sided $z$-test, while as $G$’s distribution lies further and further to the right of 0 and of $F$’s distribution, the re-centered test’s asymptotic power approaches that of the one-sided $z$-test. Although $G$’s distribution depends on many quantities, its mean is inversely proportional to $\tau_{\lambda}$, suggesting that strong LASSO performance moves $G$’s distribution to the right, matching the intuition from Section 2.2 that the power benefits of our re-centering increase with the performance of the LASSO. In Appendix E we study further how $G$’s distribution depends on the data-generating distribution and how this translates to power for the re-centered and $\ell$-tests. Key takeaways include that the match between the asymptotic theoretical power from Theorem 3 and empirical finite-sample power is quite strong, and that the asymptotic benefit of re-centering in Setting 1 is quite resilient to coefficient density, providing considerable gains across signal sizes even when $B_{0}$ is a Gaussian distribution (Figure 9) and showing no apparent power loss to the standard 2-sided test even in the “densest” setting when all entries of $\bm{\beta}$ take the same non-zero value (Figure 8). Note we can think of Gaussian coefficients as somewhat/approximately sparse, in the sense that some coefficients are closer to zero than others, even if none are exactly zero; this seems to be sufficient for the $\ell$-test to gain power via re-centering. Section 6 contains further empirical demonstrations and discussion of the power benefits of re-centering.
Remark 1.
Although the mathematically simpler direct re-centering approach we consider in this section leads to a viable alternative to the $\ell$-test, the focus of this paper remains on the $\ell$-test as presented in Section 2 because it (a) is more interpretable for most (especially non-statistician) users as it is based on a natural estimator of $\beta_{j}$ under sparsity, (b) has similar run-time as the re-centering test (both being dominated by the cross-validated LASSO), and (c) has an immediate extension to post-selection inference (Section 5).
@@ 4 $\ell$-test confidence intervals
If we can use the $\ell$-test to test $H_{j}(\gamma):\beta_{j}=\gamma$ for any $\gamma\in\mathbb{R}$, then this family of tests can be inverted to obtain a valid confidence region for $\beta_{j}$. But extending the $\ell$-test to $H_{j}(\gamma)$ is straightforward, since $\bm{y}$ satisfying $H_{j}(\gamma)$ is equivalent to $\bm{y}-\gamma\bm{X}_{j}$ satisfying $H_{j}(0)=H_{j}$, so we can simply apply the regular $\ell$-test (exactly as detailed in Section 2) to the data $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$. Defining $p_{j}^{\hat{\lambda}_{j}(\gamma)}(\gamma)$ as the $\ell$-test $p$-value for $H_{j}(\gamma)$, the $100(1-\alpha)\%$ $\ell$-test confidence region is given by $\{\gamma\in\mathbb{R}:p_{j}^{\hat{\lambda}_{j}(\gamma)}(\gamma)>\alpha\}$, and for interpretability purposes, we take its convex hull (i.e., the smallest interval containing it) as our $\ell$-test confidence interval:
@@
Computationally, one may be concerned that computing $\hat{C}_{j}$ requires many LASSO runs for a fine grid of $\gamma$ values. It is known (lars, piecewiselinear) that the LASSO solution is piecewise linear in $\lambda$ and that these paths can be generated by efficient algorithms, but in Appendix F, we show that the LASSO solution ($\hat{\bm{\beta}}^{\lambda}_{-j}(\gamma)$) is also piecewise linear in $\gamma$ (for fixed $\lambda$) and we provide an algorithm to efficiently generate these paths as well. Combining these two path-generating algorithms (in $\lambda$ and $\gamma$) provides an efficient way to share computation to efficiently compute all the LASSO solutions needed for $\hat{C}_{j}$.
@@ 5 $\ell$-test inference conditional on LASSO selection
The $\ell$-test $p$-value’s distributional form (2.7) makes it extremely straightforward (both conceptually and computationally) to adjust it to be conditionally valid given $\hat{\beta}_{j}^{\lambda}\neq 0$: simply divide the $\ell$-test $p$-value $p_{j}^{\lambda}$ by $r_{j}^{\lambda}:=\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}\neq 0\mid\bm{S}^{(j)})=\mathbb{P}_{H_{j}}(u_{1}\notin[v_{-},v_{+}]\mid\bm{S}^{(j)})=1-F_{u}(v_{+})+F_{u}(v_{-})$ to get $\underline{p}_{j}^{\lambda}:=p_{j}^{\lambda}/r_{j}^{\lambda}$. The fact that, under $H_{j}$, $\underline{p}_{j}^{\lambda}$ has a $\mathrm{Unif}(0,1)$ distribution conditional on $\bm{S}^{(j)}$ and $\hat{\beta}_{j}^{\lambda}\neq 0$ follows because $r_{j}^{\lambda}$ is exactly the supremum value $p_{j}^{\lambda}$ can take as long as $\hat{\beta}_{j}^{\lambda}\neq 0$, and the density of $p_{j}^{\lambda}$ between 0 and $r_{j}^{\lambda}$ is uniformly distributed (see Section 2.3); it follows further that $\underline{p}_{j}^{\lambda}$’s null distribution conditional only on $\hat{\beta}_{j}^{\lambda}\neq 0$ is also $\mathrm{Unif}(0,1)$.
Corollary 2 (Validity of the conditional $\ell$-test).
@@
Now for obtaining post-LASSO-selection valid confidence interval for $\beta_{j}$, we need to invert a conditionally valid test for $H_{j}(\gamma),\gamma\in\mathbb{R}$. For testing $H_{j}(\gamma)$, following the suggestion in Section 4, the test statistic should be based on the LASSO estimate on $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$, that is $|\hat{\beta}_{j}^{\lambda}(\gamma)|$, whereas the model selection event is still based on the original, un-centerd LASSO estimate, $\hat{\beta}_{j}^{\lambda}$. Furthermore, one can choose to use a different penalty parameter for the test statistic than the one used for the selection event. This prompts us to understand tests for $H_{j}(\gamma)$ based on $|\hat{\beta}_{j}^{\lambda_{\ell}}(\gamma)|$, valid conditionally on $\{\hat{\beta}_{j}^{\lambda_{s}}\neq 0\}$, where $\lambda_{\ell}$ and $\lambda_{s}$ need not be the same. Because now the test-statistic and the selection event are based on different LASSO estimates, a conditional $p$-value would not have such a simple form as for $\underline{p}^{\lambda}_{j}$, but we can still obtain a valid $p$-value using CDF transforms if we can characterize the distribution of $|\hat{\beta}_{j}^{\lambda_{\ell}}(\gamma)|\mid\{\hat{\beta}_{j}^{\lambda_{s}}\neq 0\}$, under $H_{j}(\gamma)$.
First note that, as discussed in Section 4, Lemma 1 can be applied to $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$ under $H_{j}(\gamma)$ to show that $\bm{S}^{(j)}(\gamma)=(\bm{X}_{j}^{T}\bm{y},\hat{\sigma}_{j}(\gamma))$ is sufficient under $H_{j}(\gamma)$, where, $\hat{\sigma}_{j}(\gamma)=\|(\bm{I}-\bm{P}_{-j})(\bm{y}-\gamma\bm{X}_{j})\|$, and that $\bm{y}-\gamma\bm{X}_{j}$ can be written as $\bm{P}_{-j}\left(\bm{y}-\gamma\bm{X}_{j}\right)+\hat{\sigma}_{j}(\gamma)\bm{V}\bm{u}^{\gamma}$, with $\bm{u}^{\gamma}\mid\bm{S}^{(j)}(\gamma)\stackrel{{\scriptstyle H_{j}(\gamma)}}{{\sim}}\mathrm{Unif}(\mathbb{S}^{n-d})$. In light of this result, one can now apply Theorem 1 to $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$ to conclude that $\hat{\beta}_{j}^{\lambda_{\ell}}(\gamma)\leq b$ if and only if $u_{1}^{\gamma}\leq\Lambda_{j}^{*}(b,\mathrm{sign}(b);\gamma)$, where $\Lambda_{j}^{*}(b,\epsilon;\gamma)$ is exactly equal to $\Lambda(b,\epsilon)$ but with $\bm{y}$ replaced with $\bm{y}-\gamma\bm{X}_{j}$ and $\mathrm{sign}(0):=1$. In fact Theorem 4 (that characterizes the distribution of $\hat{\beta}_{j}^{\lambda}\mid\bm{S}^{(j)}(\gamma)$ under $H_{j}(\gamma)$) in Appendix A shows that this same $u_{1}^{\gamma}$ can be used to characterize the event $\{\hat{\beta}_{j}^{\lambda_{s}}=0\}$, stating it is equivalent to $\{u^{\gamma}_{1}\in[\Lambda_{j}(0,\pm 1;\gamma)]\}$, where,
@@
Like liu2018powerful but unlike, e.g., lassoinference, our conditional inferences do not condition on anything about the LASSO’s selection except that it selects the $j^{\mathrm{th}}$ coefficient. As mentioned in Section 1.3, the key difference between our conditional inference and liu2018powerful’s is essentially the same as the difference between the (unconditional) $\ell$-test and the $t$-test, and indeed in Section 6.4 we find that our conditional confidence intervals improve over those of liu2018powerful similarly to how the unconditional $\ell$-test confidence intervals from Section 4 improve over standard $t$-test confidence intervals. Hence, we can understand when and how the conditional $\ell$-test outperforms liu2018powerful’s method in the same way that we understand the analogous outperformance of the $\ell$-test over the $t$-test; see Sections 2.2 and 3 for intuitive and formal details on this outperformance, respectively.